OpenStudy (anonymous):

Solve -3x^2 - 4x - 4 = 0 using the quadratic formula.

jimthompson5910 (jim_thompson5910):

Do you know the quadratic formula?

jimthompson5910 (jim_thompson5910):

If you do, how far did you get?

OpenStudy (anonymous):

would a be 3 and b be 4 and c be 4?

jimthompson5910 (jim_thompson5910):

you multiplied everything by -1 right?

jimthompson5910 (jim_thompson5910):

if so, then yes, a = 3, b = 4 and c = 4

OpenStudy (anonymous):

okay now i'm stuck at 4 \[4\pm√64 divide everything by -6\]

jimthompson5910 (jim_thompson5910):

\[\Large x = \frac{-b\pm\sqrt{b^2-4ac}}{2a}\] \[\Large x = \frac{-(4)\pm\sqrt{(4)^2-4(3)(4)}}{2(3)}\] \[\Large x = \frac{-4\pm\sqrt{16-(48)}}{6}\] \[\Large x = \frac{-4\pm\sqrt{-32}}{6}\] \[\Large x = \frac{-4+\sqrt{-32}}{6} \ \text{or} \ x = \frac{-4-\sqrt{-32}}{6}\] Making sense so far?

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

what's next?

OpenStudy (anonymous):

since the radicand is negative u have make it a positive so they will be i√32

jimthompson5910 (jim_thompson5910):

you can simplify \(\Large \sqrt{32}\) also

OpenStudy (anonymous):

and then divide both terms by 6?

jimthompson5910 (jim_thompson5910):

\[\Large \sqrt{32} = \sqrt{16*2}\] \[\Large \sqrt{32} = \sqrt{16}*\sqrt{2}\] \[\Large \sqrt{32} = 4\sqrt{2}\]

jimthompson5910 (jim_thompson5910):

so our updated steps become \[\Large x = \frac{-b\pm\sqrt{b^2-4ac}}{2a}\] \[\Large x = \frac{-(4)\pm\sqrt{(4)^2-4(3)(4)}}{2(3)}\] \[\Large x = \frac{-4\pm\sqrt{16-(48)}}{6}\] \[\Large x = \frac{-4\pm\sqrt{-32}}{6}\] \[\Large x = \frac{-4+\sqrt{-32}}{6} \ \text{or} \ x = \frac{-4-\sqrt{-32}}{6}\] \[\Large x = \frac{-4+4i\sqrt{2}}{6} \ \text{or} \ x = \frac{-4-4i\sqrt{2}}{6}\]

OpenStudy (anonymous):

will it be \[-2 \pm 4i√2 divided all by 3?\]

jimthompson5910 (jim_thompson5910):

you need to divide everything (except for the stuff in the square root) by 2, so \[\Large x = \frac{-b\pm\sqrt{b^2-4ac}}{2a}\] \[\Large x = \frac{-(4)\pm\sqrt{(4)^2-4(3)(4)}}{2(3)}\] \[\Large x = \frac{-4\pm\sqrt{16-(48)}}{6}\] \[\Large x = \frac{-4\pm\sqrt{-32}}{6}\] \[\Large x = \frac{-4+\sqrt{-32}}{6} \ \text{or} \ x = \frac{-4-\sqrt{-32}}{6}\] \[\Large x = \frac{-4+4i\sqrt{2}}{6} \ \text{or} \ x = \frac{-4-4i\sqrt{2}}{6}\] \[\Large x = \frac{-2+2i\sqrt{2}}{3} \ \text{or} \ x = \frac{-2-2i\sqrt{2}}{3}\]

jimthompson5910 (jim_thompson5910):

Now you're done. Those are the exact solutions. The approximate solutions can be found with a calculator (find the real and imaginary parts separately)

OpenStudy (anonymous):

ohhhhh ok

OpenStudy (anonymous):

thanks

jimthompson5910 (jim_thompson5910):

np